### v1

# Mitigating “Epistemic Debt” in Generative AI-Scaffolded Novice Programming using Metacognitive Scripts

Abstract.
The democratization of Large Language Models (LLMs) has given rise to “Vibe Coding”, a workflow where novice programmers prioritize semantic intent over syntactic implementation. While this lowers barriers to entry, we hypothesize that without pedagogical guardrails, it is fundamentally misaligned with cognitive skill acquisition. Drawing on Kirschner’s (2026) distinction between Cognitive Offloading and Cognitive Outsourcing, we argue that unrestricted AI encourages novices to outsource the Intrinsic Cognitive Load required for schema formation, rather than merely offloading Extraneous Load. This accumulation of “Epistemic Debt” creates “Fragile Experts” — those whose high functional utility is masked by critically low corrective competence.
To quantify and mitigate this debt, we conducted a between-subjects experiment ($N=78$) using a custom Cursor IDE plugin backed by Claude 3.5 Sonnet. Participants were recruited via Prolific and UserInterviews.com to represent “AI-Native” learners. We compared three conditions: Manual (Control), Unrestricted AI (Outsourcing), and Scaffolded AI (Offloading). The Scaffolded condition utilized a novel “Explanation Gate,” leveraging a real-time LLM-as-a-Judge framework to enforce a “Teach-Back” protocol before generated code could be integrated.
Results reveal a “Collapse of Competence”: while Unrestricted AI users matched the productivity of the Scaffolded group ($p<.001$ vs. Manual), they suffered a 77% failure rate in a subsequent AI-Blackout maintenance task, compared to only 39% in the Scaffolded group. Qualitative analysis suggests that successful vibe coders naturally engage in self-scaffolding, treating the AI as a consultant rather than a contractor. We discuss the implications for the long-term maintainability of AI-generated software and propose that future learning systems must enforce Metacognitive Friction to prevent the mass production of unmaintainable code.
Vibe Coding, Cognitive Outsourcing, Cognitive Offloading, Epistemic Debt, Script Theory of Guidance, Metacognitive Friction, Generative AI, Large Language Models, Computer Science Education, Software Engineering, Conversational Agents, Intelligent Tutoring Systems
copyright: rightsretainedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: The 13th ACM Conference on Learning at Scale; June 2026; Location TBDisbn: 978-1-4503-XXXX-X/26/06ccs: Social and professional topics Computing educationccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Human computer interaction (HCI)
1. Introduction
A prevailing narrative in the technology sector suggests that software engineering as a discipline is nearing its end. Industry leaders and commentators have argued that with the advent of advanced Generative AI (GenAI), “English is the new programming language” (World Governments Summit, 2024; Karpathy, 2023), implying that the barrier to creating software has effectively dissolved. The assumption is that if AI can generate functional code from high-level natural language prompts—a practice colloquially known as “Vibe Coding” (Karpathy, 2024)—then the necessity for human expertise in syntax and logic is obsolete. Indeed, the long march of software engineering has moved toward abstracting away low-level details—from assembly programming to providing expressive power through language definitions closer to human natural language, such as Python. In that trajectory, treating human natural language as the next programming abstraction seems inevitable. However, from the standpoint of learning and the long-term maintainability of software, we posit a contrasting view.
While GenAI has undoubtedly democratized the creation of software, we hypothesize that unrestricted reliance on these tools, particularly in educational contexts, is engineering a latent crisis. If the next generation of developers learns to orchestrate “vibes” (intent) without understanding the underlying mechanics, we are not witnessing the democratization of engineering, but the accumulation of Epistemic Debt.
Epistemic Debt describes the accumulation of functional software artifacts that the user “owns” legally but does not “own” cognitively. Unlike technical debt, which resides in the codebase, epistemic debt resides in the mind of the developer. When the inevitable abstraction leak occurs (Spolsky, 2002)—whether through API deprecation, hallucinated vulnerabilities, or scaling failures—the “Vibe Coder” lacks the mental model required to intervene. If this debt is not serviced during the learning process, the field faces a risk of a “Collapse of Competence,” where the workforce is replete with “Fragile Experts” who are skilled at adopting tools (high utility) but incapable of maintaining the systems they build (low comprehension).
Early hypotheses and evidence observing the over-reliance on GenAI in educational contexts have already started to question whether GenAI acts as a tool or is becoming a crutch (Balzan et al., 2026). According to a CDT national poll, 70% of teachers worry AI might weaken critical thinking and research skills, and more than half of students say they feel less connected to teachers when using AI tools (Laird et al., 2025). Surveys regarding ChatGPT use just a year after its unveiling already raised fears that “learners may rely too heavily on the model” (Kasneci et al., 2023). Furthermore, “diminished epistemic vigilance, superficial learning and emotional dependence on AI” (Yan et al., 2025), as well as metacognitive laziness (Fan et al., 2025), have been observed in specific studies. Indeed, students themselves report cognitive offloading and over-reliance on GenAI tools (Wei et al., 2025). A useful metaphor to understand this phenomenon was posited by Kirschner (2026): the distinction between cognitive offloading and cognitive outsourcing, founded in decades of cognitive and learning science research. On that basis, it is already possible to posit that our hypothesis, pointing to the accumulation of epistemic debt from the over-reliance on GenAI tools and the resulting competence collapse, will hold. Consequently, we must go a step further and ask whether this breakdown is inevitable.
Can we design AI interactions and interventions that allow novices to leverage the speed and productivity benefits of GenAI tools without sacrificing the depth of their learning? In other words, can we arrive at a best-of-both-worlds scenario where high functional utility is achieved while learners retain enough expertise to produce comparable corrective competence?
To that end, we once again turn to learning science research, drawing inspiration from Teachable Agents (Biswas et al., 2005) to design an “Explanation Gate” intervention. This intervention requires novice learners, in a scaffolded setting, to teach-back what they have learned and understood to their AI chatbot before being allowed to use the generated code. We hypothesize that this introduction of “Metacognitive Friction” can restore learner comprehension without significantly degrading productivity. Our research questions are as follows:
Our research questions are as follows:
  • RQ1 (Functional Utility & Velocity): To what extent does the introduction of automated metacognitive friction (the “Explanation Gate”) impact the immediate functional utility and development velocity of novice programmers compared to unrestricted GenAI usage?
  • RQ2 (Corrective Competence & Epistemic Debt): Does unrestricted GenAI access lead to a collapse of competence in subsequent maintenance tasks, and can an explanation gate scaffold mitigate this accumulation of epistemic debt?
  • RQ3 (Interactional Stances): What qualitative differences in human-AI interaction strategies (e.g., “Consultant” vs. “Contractor” stances) correlate with the retention of corrective competence in unrestricted GenAI environments?
Our contributions are as follows:
  1. (1)
    Theoretical Operationalization: We define and operationalize the transition from Cognitive Offloading to Cognitive Outsourcing within the specific context of GenAI-assisted programming.
  2. (2)
    Empirical Quantification of Epistemic Debt: We provide experimental evidence ($N=78$) demonstrating that while unrestricted AI preserves functional utility, it significantly degrades the corrective competence required for software maintenance.
  3. (3)
    Scalable Scaffolding Design: We introduce and evaluate the VibeCheck plugin, demonstrating that LLM-mediated metacognitive friction can restore human comprehension without significantly sacrificing AI-driven productivity gains.
2. Theoretical Framework
Our investigation is grounded in a synthesis of three complementary theories of learning: Cognitive Load Theory (specifically the distinction between Offloading and Outsourcing) (Sweller, 1988), the Script Theory of Guidance (SToG) (Fischer et al., 2013), and the “Desirable Difficulties” framework (Bjork, 1994). We argue that the interaction between novice programmers and Large Language Models (LLMs) represents a novel pedagogical context where traditional scaffolding metaphors break down, requiring a re-evaluation of how automated assistance supports, or subverts, schema acquisition.
2.1. Cognitive Load: Offloading vs. Outsourcing
A critical error in current EdTech discourse is the conflation of efficiency (productivity) with learning (schema construction). To dismantle this, we rely on Sweller’s Cognitive Load Theory (Sweller, 1988), specifically the interplay between three types of load:
  1. (1)
    Intrinsic Load: The inherent complexity of the material (e.g., control flow logic, variable state).
  2. (2)
    Extraneous Load: The mental effort imposed by the instructional format (e.g., confusing syntax, IDE configuration).
  3. (3)
    Germane Load: The mental effort dedicated to processing information into Long-Term Memory (LTM) schemas.
Expert developers use tools to minimize Extraneous Load, freeing up working memory for high-level problem solving (Germane Load). Kirschner (Kirschner, 2026) defines this as Cognitive Offloading. However, novices lack the schemas to distinguish between syntax (extraneous) and logic (intrinsic). When a novice uses GenAI to Vibe Code, the tool handles both the syntax and the underlying logic. This transition from Offloading to Cognitive Outsourcing allows the learner to bypass the Intrinsic Load entirely (Kirschner, 2026). Without engaging with Intrinsic Load, Germane processing cannot occur, and the learner accumulates epistemic debt, a state where functional output exists without the corresponding mental models to maintain it.
2.2. Overscripting
Fischer’s Script Theory of Guidance (SToG) posits that external scripts (prompts, roles, instructions) are necessary to help learners coordinate complex tasks (Fischer et al., 2013). Ideally, these scripts provide a temporary structure that is gradually “faded” as the learner’s internal scripts (schemas) mature (Wecker and Fischer, 2010).
We argue that the “Copy-Paste” AI workflow represents a pathological case of Overscripting (Dillenbourg, 2002). Because standard LLMs provide complete, high-fidelity solutions instantly, they act as a rigid script that replaces, rather than regulates, the learner’s cognitive activity. Unlike human tutors who withhold answers to prompt reflection, an unrestricted LLM resolves the problem space immediately, denying the learner the “productive failure” necessary for transfer (Kapur, 2008).
To counteract this, we introduce the concept of Metacognitive Friction. This design paradigm deliberately slows down the interaction loop, forcing the learner to engage in generative processing. This aligns with Bjork’s concept of Desirable Difficulties (Bjork, 1994), specifically the “Generation Effect,” which posits that long-term retention is improved when learners must actively retrieve and reformulate knowledge rather than passively consuming it.
2.3. Scalable Assessment as the Gatekeeper
Historically, the Teachable Agent paradigm, where students learn by teaching a computer agent, has been shown to be highly effective for deep learning (Biswas et al., 2005). However, verifying that a student has truly “taught” the agent (rather than gaming the system (Baker, 2005)) traditionally required human Teaching Assistants (TAs), limiting the scalability of such interventions.
Recent advancements in LLM-Based Assessment allow us to automate this verification at scale. By leveraging an “LLM-as-a-Judge” framework (Choma et al., 2025), we can convert the Teachable Agent model into a real-time “Explanation Gate.” This system acts as a semantic validity check, ensuring that the learner’s vibe (intent) is backed by sufficient syntactic understanding before code integration is permitted. This effectively closes the loop on Outsourcing, forcing the learner to convert the AI’s output back into their own internal representation before moving forward.
3. Methodology
3.1. Participants & Recruitment
We recruited $N=78$ participants located in the United States for a 2-hour synchronous remote session. To ensure a representative sample of AI-Native learners, we utilized a dual-channel sourcing strategy to capture both traditional students and vocational learners:
  1. (1)
    Prolific ($n=53$): Targeting current undergraduate students majoring in Computer Science or related STEM fields.
  2. (2)
    UserInterviews.com ($n=25$): Targeting recent graduates (<12<12 months) of intensive coding bootcamps to capture the vocational segment.
Screening Criteria: Participants were screened for basic familiarity with JavaScript (i.e., they could write a loop and define a function) but no professional experience with React.js (the target framework).
Demographics: The sample was 34% female, 64% male, and 2% non-binary. While we attempted to stratify for gender balance, the pool reflected the broader demographics of the sourcing platforms and the domain. The mean age was 22.1 years ($SD=1.8$, Range=19–29). All participants were compensated at a rate of \$15 USD/hour.
3.2. The Coding Environment
We developed a custom experimental environment to mimic a real-world Vibe Coding setup while maintaining experimental control.
  • IDE: All participants used Cursor IDE, a fork of VS Code that integrates LLM features natively.
  • Generator Model: We utilized Claude 3.5 Sonnet (via API) for all code generation, chosen for its dominance in coding benchmarks as of the time this study was conducted in Fall 2025 (Anthropic, 2024).
  • Plugin Architecture: We developed a custom VS Code extension, VibeCheck, which acted as a middleman between the LLM Sidebar and the editor. As illustrated in Figure 1, the extension implemented two core logic layers:
    1. (1)
      Rule Enforcement Engine: A background listener that intercepted input events from two sources: user keystrokes/clipboard and the LLM Sidebar’s “Apply” signal. It routed actions based on the experimental condition—blocking external pastes for the Manual group, or intercepting the merge signal for the Scaffolded group.
    2. (2)
      Explanation Gate Manager (The Friction Loop): A state machine responsible for the intervention. Upon receiving a halt signal from the Enforcement Engine, this manager rendered a blocking modal overlay. It orchestrated an asynchronous loop with the Judge API, preventing the code merge until the Judge returned a “Pass” evaluation, effectively closing the feedback loop shown in the diagram.
    3. (3)
      Telemetry Logger: Recorded every keystroke, chat interaction, and state change with millisecond-precision timestamps for post-hoc analysis.
System architecture flowchart showing input from Sidebar and Keystrokes being processed by Rule Enforcement Engine and potentially gated by the Explanation Gate Manager. Refer to caption
Figure 1. System Architecture of the VibeCheck Plugin. The diagram illustrates how the Rule Enforcement Engine intercepts the “Apply” signal from the LLM Sidebar and conditionally routes it through the Explanation Gate Manager, enforcing a Metacognitive Friction loop via the external Judge API.
3.2.1. The “Judge” Agent
To enable scalable assessment, the Explanation Gate utilized a secondary LLM agent to evaluate student explanations in real-time. We selected GPT-4o (OpenAI) as the judge to minimize self-preference bias (where a model prefers its own output style) (Wataoka et al., 2024; Chen et al., 2025). The judge was configured with a temperature of 0.1 to ensure deterministic grading.
System Prompt: The judge was initialized with the following rubric based on the SOLO Taxonomy (Biggs and Collis, 1989):
You are a Teaching Assistant for a React course. Your goal is to evaluate if the student truly understands the code they are trying to merge.
Context: The student has generated code using AI and must explain it to you.
Rubric: Grade their explanation on a scale of 1-5:
  • 1 (Pre-structural): Tautological (e.g., It adds the button).
  • 3 (Relational): Explains how components interact (e.g., It uses the state hook to update the list).
  • 5 (Extended Abstract): Explains edge cases or architectural impact.
Output: JSON format { "score": number, "feedback": string }. If score ¡ 3, provide specific, Socratic feedback to guide them (e.g., You mentioned state, but how does the useEffect dependency array affect the update loop?). Do not reveal the answer.
The student saw the modal prompt: “Wait! Before applying this code, explain its causal logic. How does it handle state updates?” If the Judge returned a score below 3, the modal remained locked, displaying the Judge’s feedback. This cycle repeated until a passing score was achieved.
The use of the SOLO taxonomy for evaluating student explanations was derived with the help of an expert learning designer as the human-in-the-loop using the process described in Choma et al. (2025).
Table 1. SOLO Taxonomy Calibration for Automated Judge Evaluation
Score SOLO Level Student Explanation Example
1 Pre-structural “I added a button so that when you click it the course is added to the list and it updates.”
2 Unistructural “It uses the enrollCourse function to send the ID to the database and then changes the UI.”
3 Relational “The await keyword ensures the app waits for the DB confirmation before updating the local courses state, preventing UI desync.”
5 Ext. Abstract “It implements optimistic updates by updating the state first but uses a try/catch to rollback the UI if the fetch request fails, ensuring eventual consistency.”
3.3. Experimental Conditions
Participants were assigned via a Stratified Randomization protocol. Within each recruitment channel, participants were randomly assigned to one of three conditions to ensure equal distribution ($n=26$ per group).
3.3.1. Group A: Manual (Control)
Participants used the standard VS Code environment. They had access to official React documentation but were strictly forbidden from using Generative AI tools.
Enforcement: The VibeCheck Rule Enforcement Engine monitored clipboard events; paste events exceeding 20 characters from external sources triggered a warning flag. Additionally, screen recordings of a random 10% sample were manually reviewed for compliance.
3.3.2. Group B: Unrestricted AI (Outsourcing)
Participants used Cursor IDE with free access to Claude 3.5 Sonnet in the sidebar. The Apply button was enabled instantly. This feature automatically diffs and merges generated code into the active file, mimicking the default Vibe Coding experience.
3.3.3. Group C: Scaffolded AI (Offloading)
Participants used Cursor IDE with the extension active. The Apply button was disabled by default. To unlock it, the Explanation Gate Manager triggered a modal asking: “Explain the causal logic of this block. How does it handle state updates?”
The Judge: The student’s response was sent to a background GPT-4o agent. We purposefully selected a model family different from the generator (Claude) to minimize self-preference bias in evaluation, as explained in the previous section.
Feedback Loop: If rejected, the student received the feedback (e.g., “You identified what variables are used, but not why the effect hook is needed here”) and had to retry.
3.4. Task Protocol
The experiment consisted of two distinct phases designed to evaluate the trade-off between the immediate production of software and the long-term acquisition of learner competence.
3.4.1. Phase 1: Functional Utility
Over a 90-minute period, participants were tasked with building a “Student Course Scheduler” React application. This phase was designed to measure Functional Utility. The application requirements included:
  1. (1)
    Authentication: A functional mock login interface.
  2. (2)
    Data Rendering: Displaying course metadata from a provided JSON source.
  3. (3)
    Stateful Enrollment: Implementing a selection system to add courses.
  4. (4)
    Conflict Logic: Algorithmic detection to prevent overlapping schedules.
During this phase, the Rule Enforcement Engine and Explanation Gate Manager were active. For Group C, every attempt to integrate logic triggered the Metacognitive Friction loop, requiring a passing explanation before code could be merged.
Primary Metric: The Functional Utility Score (0–100), calculated via a headless Puppeteer/Jest test harness that executed 12 end-to-end assertions. Each assertion was weighted equally (8.33 points).
Table 2. 12 Component Assertions of the Functional Utility Score for a Total of 100 Points
ID Assertion Description
A-01 User can log in using mock credentials and reach the dashboard.
A-02 Course catalog successfully fetches and renders 20+ items from JSON.
A-03 Course list displays correct metadata (Instructor, Credits, Time).
A-04 Clicking “Enroll” adds a course to the user’s sidebar state.
A-05 Duplicate enrollment in the same course is blocked by state logic.
A-06 The system identifies time-overlaps between two distinct courses.
A-07 A warning modal appears when a conflict is detected.
A-08 Total credit count updates dynamically as courses are added/removed.
A-09 Enrollment persistence: data remains after simulated page refresh.
A-10 Search functionality filters the course catalog by name/instructor.
A-11 Conflict detection logic correctly handles back-to-back classes.
A-12 Final schedule export generates a valid summary string.
3.4.2. Phase 2: Corrective Competence
Immediately following Phase 1, we simulated an “abstraction leak” to measure the participants’ Corrective Competence. AI access was revoked for all groups, and we used a git patch to inject a Logic Bomb into their functioning codebase.
We introduced a race condition in the enrollCourse function by stripping the await keywords from asynchronous database calls and deleting the optimistic UI rollback logic. Under simulated network latency, the UI would optimistically show a course as added, but the underlying state would fail to sync, causing the course to vanish upon browser refresh (a “ghost course”). Participants were given the final 30 minutes to identify the cause of the ghost courses and manually implement a fix without AI assistance.
It is important to distinguish between general debugging and the measurement of Corrective Competence. To ensure we were measuring the latter, the Logic Bomb was not a novel bug, but a structural regression of the participant’s own existing code. Because all participants in Groups B and C had successfully implemented a functional enrollment system in Phase 1 (verified by the test harness), they had all previously “owned” a version of the code where the asynchronous await and rollback logic were present. Thus, the Phase 2 task did not require the discovery of a new architectural pattern, but rather the recognition and restoration of a logic flow they had ostensibly integrated already prior. Failure to resolve this regression served as a direct proxy for the Epistemic Debt accumulated during the initial construction.
Primary Metric: Corrective Competence, measured as a binary Repair Success (resolved vs. unresolved) and Time-to-Fix.
Table 3. The Phase 2 “Logic Bomb” Regression
Original (Functional) Injected Regression (Bug)
const res = await fetch(’/enroll’); const res = fetch(’/enroll’);
if (res.ok) { updateUI(); } updateUI(); // Optimistic
else { rollback(); } // No rollback logic
4. Results
Our analysis evaluates the experimental data through the lens of the trade-off between immediate software production (Utility) and the acquisition of long-term learner competence.
4.1. RQ1: Functional Utility and Velocity
To evaluate whether the Metacognitive Friction of the Explanation Gate significantly impeded immediate software production, we analyzed both the Phase 1 Functional Utility Scores and the total time spent on the construction task.
4.1.1. Functional Utility
We verified the assumptions for parametric testing; normality was confirmed using a Shapiro-Wilk test ($W=0.98,p=.42$), and homogeneity of variance via Levene’s test ($F(2,75)=1.12,p=.33$). A one-way ANOVA revealed a highly significant effect of condition on utility ($F(2,75)=48.2,p<.001,\eta^{2}=0.56$).
As detailed in Table 4, both AI-assisted groups achieved significantly higher utility than the manual control. Post-hoc Tukey HSD tests revealed no significant difference in functional utility between the Unrestricted and Scaffolded groups ($p=.64$).
Table 4. Phase 1: Functional Utility Scores
Condition Mean Utility Score SD
Manual (Group A) 65.2% 14.3
Unrestricted AI (Group B) 92.4% 8.1
Scaffolded AI (Group C) 89.1% 9.5
4.1.2. Development Velocity and Cognitive Allocation
To evaluate the “velocity cost” of metacognitive friction, we compared total time-on-task and the allocation of Active Thinking time. We define Active Thinking as the duration spent in manual construction, code comprehension, and, for Group C, the duration of Explanation Gate interventions.
As shown in Table 5, the progress metric reveals that the Manual group exhausted the 90-minute limit having only satisfied 42% of the requirements. In contrast, both AI-enabled groups achieved 100% completion.
While the Scaffolded group encountered a significant velocity penalty compared to the Unrestricted group ($p<.001$), their “Net Velocity” remained superior to the Manual control. The friction introduced by the Explanation Gate (Median = 14.2 minutes) was effectively subsidized by the generational speed of the LLM.
Table 5. Phase 1: Time on Task and Cognitive Allocation
Condition Total Time Mean ($SD$) Active Thinking Mean ($SD$) Progress (%)
Manual (A) 88.4 (2.1) 88.4 (2.1) 42%
Unrestricted (B) 48.2 (10.4) 18.5 (4.2) 100%
Scaffolded (C) 64.6 (11.2) 36.3 (5.8) 100%
4.2. RQ2: Corrective Competence and Epistemic Debt
In Phase 2, we measured the Corrective Competence of participants after AI access was revoked. Repair Success rates for the race-condition logic bomb were analyzed using a Chi-square test of independence, yielding significant results ($\chi^{2}(2)=13.8,p=.001,V=0.42$).
  • Manual (Group A): 18/26 succeeded (69.2%). Despite lower initial utility, these participants possessed high cognitive ownership.
  • Unrestricted AI (Group B): 6/26 succeeded (23.1%). This group exhibited a clear Collapse of Competence.
  • Scaffolded AI (Group C): 16/26 succeeded (61.5%). The “Explanation-Gate” successfully mitigated the accumulation of epistemic debt.
4.3. RQ3: Interactional Stances
We investigated the 6 participants in Group B who successfully fixed the bug. We identified two distinct behaviors:
  1. (1)
    The Contractor Stance: Prompts focused on functional output.
  2. (2)
    The Consultant Stance: Successful outliers consistently prompted for clarification.
Successful participants utilized a mean of 4.2 explanatory prompts per session, compared to 0.4 for those who failed ($t(24)=5.1,p<.001$).
Table 6. Representative Individual Results showing the divergence in Phase 2.
Table showing Phase 1 utility scores against Phase 2 repair results and time to fix across different conditions and participant IDs.
ID Condition Utility (Ph 1) Repair Result (Ph 2) Fix Time Dominant Stance (Ph 1)
P-12 Manual 72% Success 11m 45s Manual Reasoning
P-23 Unrestricted 95% Fail Timeout Contractor (Directive)
P-35 Unrestricted 98% Success 21m 10s Consultant (Explanatory)
P-48 Scaffolded 90% Success 13m 30s Enforced Consultant
P-55 Scaffolded 82% Fail Timeout Tautological (Low SOLO)
4.4. The Consultant vs. The Contractor
Our analysis of the interaction logs in Group B (Unrestricted) reveals a critical behavioral predictor of corrective competence. We identified two dominant interactional stances:
The Contractor Stance::
These participants treated the LLM as a low-level worker. Prompts were purely imperative (e.g., “Make the button blue,” “Add the filter”). Telemetry shows these users rarely scrolled through the generated code before clicking “Apply.” This stance maximize immediate velocity but results in 100% Cognitive Outsourcing.
The Consultant Stance::
Outliers who succeeded in Phase 2 treated the LLM as a senior partner. Their prompts were interrogative (e.g., “Why did you use a useEffect here?” or “Can you explain this state change?”). These users engaged in Self-Scaffolding, effectively “paying” their epistemic debt voluntarily.
The VibeCheck plugin effectively forces Contractors to behave like Consultants. This suggests that the future of AI-assisted learning is not just about the quality of the AI’s code, but about designing interfaces that preclude the Contractor stance.
4.5. Qualitative Perceptions of Friction
Post-session surveys revealed a “Bimodal Sentiment” toward the Explanation Gate. While 72% of Group C participants initially described the gate as “annoying” or “an obstacle,” 64% of those who successfully fixed the Phase 2 logic bomb admitted that the gate was the only reason they knew where to look for the bug.
One participant (P-48) noted: “I was going to just click ’Apply’ and move on, but the modal made me realize I didn’t actually know why the state was updating. It forced me to read the code I just asked the AI to write.” This supports our hypothesis that Metacognitive Friction converts passive outsourcing into active offloading. Nevertheless, future designs should potentially balance beneficial friction with annoyance in order to improve the uptake of these affordances voluntarily and in practice.
4.6. Distribution of Explanation Quality in the Scaffolded Condition
We analyzed a total of 412 individual explanation attempts recorded across 26 participants in Group C. These attempts occurred over 172 discrete “gate encounters,” averaging 6.6 encounters per participant, a frequency that aligns with the six major architectural milestones of the task.
On average, participants required 2.4 attempts ($SD=1.1$) to satisfy the Judge API’s requirement for a SOLO Level 3 (Relational) explanation.
As shown in Figure 2, the modal peak at two attempts (n=85 encounters) illustrates the impact of the Socratic feedback. The most common failure mode (62% of rejected attempts) was the “Tautological Loop,” where students repeated the AI’s prompt or variable names without explaining causal relationships. The “Relational” breakthrough typically followed the first round of Judge feedback, effectively shifting the learner from the passive Contractor stance to an active Consultant stance.
A histogram showing the number of attempts per Explanation Gate encounter, with a clear peak at 2 attempts Refer to caption
Figure 2. A histogram showing the number of attempts per Explanation Gate encounter, with a clear peak at 2 attempts
5. Discussion
Our findings provide empirical evidence for the existence of “Epistemic Debt” in AI-augmented novice programming. By comparing immediate functional utility with subsequent corrective competence, we demonstrate that the efficiency gains of Generative AI can mask a significant collapse in learner comprehension if pedagogical guardrails are not enforced.
5.1. Mapping the Cognitive Collapse: A CLT Perspective
To understand the mechanisms behind the observed competence collapse, we mapped our Phase 1 telemetry data to the three components of Cognitive Load Theory (Sweller, 1988).
5.1.1. Extraneous Load
In the Unrestricted group (B), extraneous load, i.e., the mental effort of managing syntax and IDE state, was effectively reduced to zero. While the Explanation Gate was designed to manage Germane load, our qualitative data suggests that the Scaffolded group (C) participants initially perceived it as extraneous friction. This indicates that for AI-native learners, pedagogical interventions may be misclassified as system friction, potentially leading to the “gaming the system” behaviors observed in specific outliers (e.g., P-55).
5.1.2. Germane Load
We posit that the 38.4% success rate gap between Groups B and C is the direct result of the Explanation Gate forcing Germane Load, i.e., the effort required to build a mental schema. By requiring a “Teach-Back” response passing a SOLO level 3 threshold, we ensured that the learner could not integrate code without first converting the AI’s outsourced logic into an offloaded mental representation. This confirms that metacognitive friction is a necessary catalyst for germane processing in generative environments.
5.2. The False Floor of Competence and the Complexity Ceiling
The results from Phase 1 suggest that for novice programmers, unrestricted GenAI access creates a False Floor of Competence. Group B achieved a functional utility score significantly higher than the Manual group, despite possessing arguably less cognitive ownership of the resulting codebase. These participants effectively vibe coded their way into a degree of complexity they could not sustain.
In Phase 2, this floor collapsed. The low repair success rate of the Unrestricted group suggests that when a developer outsources the intrinsic load of schema formation (Kirschner, 2026), they hit a hard Complexity Ceiling. Scaling and debugging require an understanding of structural decomposition. Because Group B participants never internalized these schemas, they treated their own code as an opaque, foreign library.
5.3. Structural Ownership of Code Artifacts
The specific nature of the regression injected in Phase 2 (see Table 3) highlights the disconnect between visual success and structural ownership. By removing the await keyword and the rollback logic, we created a state where the application still appeared to function under ideal conditions but lacked the defensive programming required for robustness.
Participants in the Unrestricted AI group (B) were notably susceptible to this “Visual Trap.” Because the AI-generated code worked immediately in Phase 1, they never engaged with the underlying asynchronous architecture. When the logic bomb was injected, their troubleshooting logs showed a preoccupation with UI components rather than the data-fetching logic. They possessed what we define as “Surface-Level Utility”; they could confirm the button existed, but they had no mental schema for the promise-based lifecycle of the enrollment event.
In contrast, the Scaffolded group (C) was forced to explain this exact lifecycle to the Judge API during construction. By articulating the relationship between the fetch request and the state update, they performed the necessary germane processing to “buy back” their epistemic debt. When the bug was injected, these participants did not simply debug; they recognized a structural regression in a system they now cognitively owned. This suggests that metacognitive friction acts as an architectural audit, forcing the developer to move beyond the “vibe” of a working UI to the reality of the underlying implementation.
5.4. Industry Implications: Realization of Epistemic Debt
The risks identified in our laboratory setting find a direct parallel in recent industry events, validating the Epistemic Debt hypothesis at scale. The Clawdbot/Moltbot incident of early 2026 (Reuters, 2026) serves as a primary case study of the dangers of high utility coupled with low corrective competence. In that event, thousands of developers deployed personal AI agents using generated configurations they did not fully understand.
While the users successfully orchestrated the functional intent, they outsourced the implementation details, specifically the binding of insecure WebSocket ports. When these systems were scanned by malicious actors, the vibe coders were helpless to intervene. Much like the participants in our Unrestricted group, these developers were effectively Senior Architects in capability but Junior Interns in comprehension. Without the foundational mental models required for security and architectural oversight, the functional utility they achieved was fundamentally fragile.
5.5. Fragile Velocity vs. Sustainable Velocity
A common critique of metacognitive interventions is that they impede development speed. However, our data suggests this is a false dichotomy. While the Explanation Gate introduced a time penalty of approximately 14 minutes in Group C, this cost was largely subsidized by the generation speed of the LLM.
We frame this trade-off as a shift from Fragile Velocity to Sustainable Velocity. The VibeCheck intervention ensures that the velocity gained through AI is maintainable by the human developer. By enforcing the “Teach-Back” protocol, the tool ensures that the intrinsic load of the task is offloaded rather than outsourced.
5.6. Self-Scaffolding and Interactional Stances
The outlier analysis in RQ3 provides a final, critical insight: accumulation of Epistemic Debt is not solely a function of the tool, but of the user’s interactional stance. Successful participants in the Unrestricted group naturally adopted a “Consultant Stance.” This suggests that while advanced learners may possess the metacognitive discipline to treat AI as a peer, novices require structural friction to prevent them from falling into the “Contractor Stance.”
Future learning systems must, therefore, move beyond passive code generation toward active metacognitive mediation. By automating the verification of student explanations through an LLM-as-a-Judge framework (Choma et al., 2025), we can scale such interventions to meet the demands of learners in the current GenAI era.
6. Threats to Validity
6.1. Internal Validity
A primary threat is the Hawthorne Effect; participants knew they were in a monitored study, which may have artificially increased their persistence in the Scaffolded group. To mitigate this, we anonymized the Judge’s feedback to appear as standard IDE warnings rather than “test” feedback. Furthermore, the 90-minute limit for Phase 1 was chosen to simulate the time pressure of a standard technical screen, grounding the velocity results in professional reality. Nevertheless, the limitation remains, despite our mitigation strategies.
6.2. External Validity
Our study focused exclusively on novice programmers (students and bootcampers). It is possible that for expert developers, who already possess robust internal mental models, the Explanation Gate would represent purely extraneous load with no germane benefit. The “Epistemic Debt” we observed is likely a function of the novice-AI dyad, and results may not generalize to experienced programmers who utilize AI for boilerplate tasks rather than core logic. In either case, approaching the problem from first principles, i.e., defining germane, intrinsic and extrinsic loads will allow for the delineation of offloading and outsourcing for the respective populations, and designing appropriate scaffolds, as a result. We posit that our study, grounded in cognitive load theory, while not externally face valid nevertheless provides directional support for the design of scaffolds for experienced programmers as well.
6.3. Construct Validity
We used “Repair Success” as a proxy for Corrective Competence. Critics might argue that this measures debugging skill rather than understanding. However, as noted in Section 3.4.2, because the bug was a regression of the user’s own code integrated moments prior, we believe it remains a valid measure of cognitive ownership.
7. Conclusion
This study has demonstrated that the democratization of software creation via GenAI carries the latent risk of an industry-wide collapse of competence. However, this breakdown is not inevitable. By deliberately introducing Metacognitive Friction into the development loop, we can leverage the immense productivity gains of LLMs while ensuring that the humans at the helm remain capable of maintaining the complex systems they build. To prevent the mass production of unmaintainable software, we must ensure that the price of velocity is not the loss of comprehension.
8. Future Work
The findings of this study open several avenues for investigating the long-term sustainability of AI-assisted learning environments. We identify four primary areas for future research:
8.1. Adaptive Scaffolding and Instructional Fading
The current implementation of the VibeCheck Explanation Gate utilizes a static threshold (SOLO Level 3). Future work should explore adaptive friction, where the complexity of the required explanation “fades” as the system detects increasing learner competence. Drawing on the Expertise Reversal Effect (Kalyuga, 2009), a system that transitionally moves from high-friction scaffolds to frictionless utility could prevent the “Fragile Expert” syndrome while eventually granting the developer the high-velocity benefits of an unrestricted environment.
8.2. Longitudinal Learning and Mental Model Decay
Our experiment measured corrective competence immediately following a construction task. However, the long-term persistence of schemas formed under metacognitive friction remains unknown. Future longitudinal studies should investigate whether “buying back” epistemic debt through teach-back protocols leads to higher retention of architectural principles over months of professional practice, or if the “Consultant Stance” naturally erodes as users face the time-pressures of industrial production.
8.3. Collaborative Human-AI-Human Systems
While this study focused on the programmer-AI dyad, modern software engineering is a social activity. Research is needed to determine how epistemic debt impacts code reviews and collaborative debugging. If a senior developer reviews code authored by a novice “Vibe Coder,” does the lack of cognitive ownership in the junior dev create a bottleneck in the team’s collective corrective competence? Extending the Explanation Gate to include collaborative “Teach-Forward” protocols between humans could mitigate these social bottlenecks.
8.4. Equity, Diversity, and the AI Experience Gap
A critical area for future inquiry is the impact of Generative AI on underrepresented groups in Computer Science. Currently, Black and Hispanic students earn close to 20% of CS degrees in the United States, combined, (Siek and Batten, 2025; National Center for Education Statistics, 2023) yet they often report higher rates of “imposter syndrome” and lower access to early-career mentorship. Generative AI has the potential to act as a “leveling agent”; however, if unrestricted AI usage leads to a disproportionate accumulation of epistemic debt among learners with less existing social capital, it may inadvertently widen the “competence gap” during technical interviews. Future research must quantify whether scaffolded AI tools like VibeCheck provide a more equitable pathway to senior-level architectural understanding for students from diverse socio-economic and racial backgrounds.